OpenStudy (anonymous):

I need help, What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?.. I got the slope but I am stuck from there, I think the slope is 2

OpenStudy (anonymous):

well the equation to get the slope is y2-y1/x2-x1

OpenStudy (anonymous):

and to get the point slope form the equation is y – y1 = m(x – x1)

OpenStudy (anonymous):

so just plug in the points into the equation

OpenStudy (anonymous):

then im really lost

OpenStudy (anonymous):

where are you lost?

OpenStudy (anonymous):

ok so the equation would be something like y=-2(x+2) ?

OpenStudy (anonymous):

(–2, 0) and (2, 8) and the equation has to be y -y1 = m(x - x1) but the coordinates are like (x1,y1) = (2,8)

OpenStudy (anonymous):

do get it a little?

OpenStudy (anonymous):

yeah im understanding a little more

OpenStudy (anonymous):

but first of all you need to know how to get the slope first do you know the slopes equation??

OpenStudy (anonymous):

i mean how to get the slope?

OpenStudy (anonymous):

yes thats y2-y1/x2-x1 right

OpenStudy (anonymous):

yes and do you know how that works?

OpenStudy (anonymous):

yes Ill get a fraction and then i simplify ?

OpenStudy (anonymous):

maybe but how did you get 2 in the first place

OpenStudy (anonymous):

I got 8/4 somehow and then I dont know what i did

OpenStudy (anonymous):

well do you know what y1 and y2 and x1 and x2 is?

OpenStudy (anonymous):

the coordinates?

OpenStudy (anonymous):

yes and specifically do you know which one it is from this is y1 and y2 and x1 and x2? (–2, 0) and (2, 8)

OpenStudy (anonymous):

0=y1 8=y2 and -2=x1 2=x2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so put it in the equation

OpenStudy (anonymous):

8-0/2-(-2) =2/4

OpenStudy (anonymous):

8/4 = 2 is m (slope)

OpenStudy (anonymous):

so then it would be y-8=2(x-2) y=2x-6?

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

y = 2x - 6 not quite 4-8

OpenStudy (anonymous):

y = 2x - 4

OpenStudy (anonymous):

but it says it wants it at point slope form just leave it as y-8=2(x-2)

OpenStudy (anonymous):

I got that the second time but it doesnt even have that as a choice of an answer.... it has y=2(x-2)

OpenStudy (anonymous):

oh then it would be the first coordinate that you must put into the equation.

OpenStudy (anonymous):

y - 0 = 2(x + 2) or y = 2(x + 2)

OpenStudy (anonymous):

okay thank you so much :)

OpenStudy (anonymous):

yw